OpenStudy (anonymous):

solve the equation as an equation reducible to the equation. (2x-4)^2-4(2x-4)-5=0

OpenStudy (anonymous):

so 2.5?

OpenStudy (aum):

(2x-4)^2 - 4(2x-4) - 5 = 0 Let t = (2x-4) t^2 - 4t - 5 = 0 Solve the quadratic for t. Then replace t with 2x-4 and solve for x.

OpenStudy (anonymous):

dont quite get it. could you brake it down for me.

OpenStudy (aum):

(2x-4)^2 - 4(2x-4) - 5 = 0 Let t = (2x-4). Replace (2x-4) by t in the above equation: t^2 - 4t - 5 = 0 Solve for t. First factor the above quadratic equation: t^2 - 4t - 5 = (t - 5)(t +1) = 0 t = 5 or t = -1 put back 2x - 4 in the place of t: 2x - 4 = 5 or 2x - 4 = -1 2x = 9 or 2x = 3 x = 9/2 or x = 3/2